### v1

Abstract
We construct a family of quantum low-density parity-check codes locally equivalent to higher-dimensional quantum hypergraph-product (QHP) codes. Similarly to QHP codes, the proposed codes have highly redundant sets of low-weight stabilizer generators, which improves decoding accuracy in a fault-tolerant regime and gives them single-shot properties. The advantage of the new construction is that it gives shorter codes. We derive simple expressions for the dimension of the proposed codes in two important special cases, give bounds on the distances, and explicitly construct some relatively short codes. Circuit simulations for codes locally equivalent to 4-dimensional toric codes show a (pseudo)threshold close to $1.1\%$, better than for toric or surface codes with a similar noise model.
I Introduction.
Quantum error correction (QEC) is one of the key enabling technologies for scalable quantum computation. Even so, despite a lot of effort in the community, it took over quarter century after discovery of QEC[1, 2] for quantum memory experiments based on surface codes[3, 4, 5] to demonstrate error rates that are well in the scalable regime[6, 7, 8, 9]. Major difficulty here is the need to operate fault-tolerantly, in the presence of errors that occur during the syndrome measurement.
A general approach to fault-tolerance (FT) is to use redundant syndrome measurements. For quantum memory and protected Clifford gates, the needed redundancy can be gained from repeated measurements, as it is sufficient just to track the Pauli basis associated with the current error. Here, the classical syndrome-based decoding problem can be solved sequentially, using measured syndrome data from a few cycles at a time[5, 10].
On the other hand, non-Clifford gates needed for universal quantum computation require real-time decoding. Even though some delay can be tolerated at the cost of increased circuit complexity, shorter lag times are strongly preferable[11]. From this viewpoint, optimal is single-shot decoding[12, 13, 14], where syndrome data from each measurement cycle is processed right after it becomes available.
Single-shot fault-tolerant QEC requires measuring some redundant stabilizer generators. This redundancy helps to control measurement errors and thus enables independent decoding of each batch of syndrome data. Such ideas go back to the seminal work[5] of Dennis et al.; various coding theory aspects have been studied in, e.g., Refs. 15, 16, 17. Single-shot properties have been studied numerically and analytically for a number of code families, including higher-dimensional hyperbolic codes[18], as well as higher-dimensional toric[19], and more general quantum hypergraph-product (QHP) codes[20, 21, 22, 23].
Unfortunately, the product constructions tend to give rather long codes. Recently, the effect of natural redundancy in two-block and generalized-bicycle (GB) codes[24] with stabilizer generators of weight six have been studied by three of the present authors[25]. These, and more general 2BGA codes[26, 27, 28], including the bivariate-bicycle (BB) codes[29, 30, 31, 32], have come to the attention as the family of codes which contains short codes with relatively high rates and distances (substantially improving over the surface/toric codes) and at the same time with relatively high thresholds. While GB codes constructed in Ref. 25 saturate an upper bound on the syndrome distance, their single-shot properties turned out to be limited by the lack of confinement[22].
More recently, Aasen et al. proposed a family of four-dimensional geometric toric codes[33, 34], locally equivalent to 4D toric codes but on a 4-torus with non-collinear periodicity vectors, analyzed their single-shot properties and performance as quantum memory, and designed a number of protected operations on logical qubits, including state injection and the entire logical Clifford group.
In this work we construct a family of multi-block chain (MBC) complexes, and their implementation using abelian-group Hoph algebras, abelian multi-cycle (AMC) codes. This construction is strictly more general than 4D geometric toric codes in Refs. 33, 34, in the sense that all their codes are included in our construction. Our MBC codes are related to higher-dimensional QHP codes[20, 21] in exactly the same way as two-block codes are related to the original (two-dimensional) QHP codes[35], and they share many properties with both the two-block codes and higher-dimensional QHP codes. In addition to the general construction, we give analytic expressions for the code dimensions in two important general cases, and also give lower (existence) and upper bounds for the distances. As an important example, we study the family of AMC codes locally equivalent to 4-dimensional toric codes. All these codes have stabilizer generators of weight $6$ and encode $k=6$ logical qubits. We explicitly construct a number of such codes, calculate their parameters, and perform circuit simulations to compute a (pseudo)threshold and analyze the accuracy of few-shot sliding-window decoding.
II Definitions.
For a given finite (Galois) field $F\equiv\mathbb{F}_{q}$, of size $|F|=q=p^{m}$ and prime characteristic $p$, an $F$-linear code of length $n$ and dimension $k$ is a linear space ${\cal C}\subseteq F^{n}$ of dimension $k$. Any set of linearly-independent rows of a generator matrix $G$ forms a basis of the code ${\cal C}\equiv{\cal C}_{G}$ generated by $G$; the code dimension is $k=\mathop{\rm rank}G$. The dual code ${\cal C}^{\perp}\equiv{\cal C}_{G}^{\perp}$ of dimension $n-k$ is formed by all vectors in $F^{n}$ orthogonal to non-zero vectors in ${\cal C}$ or, equivalently, to the rows of $G$, a check matrix of the code ${\cal C}_{G}^{\perp}$.
For a given finite field $F$ and a finite group $G$ of order $|G|=\ell$, the group Hopf algebra (also, a ring) $F[G]$ is defined as an $F$-linear space of all formal sums
$x\equiv\sum_{g\in G}x_{g}g,\quad x_{g}\in F,$ (1)
where group elements $g\in G$ serve as basis vectors, equipped with the product naturally associated with the group operation,
$ab=\sum_{g\in G}\biggl{(}\sum_{h\in G}a_{h}b_{h^{-1}g}\biggr{)}g,\quad a,b\in F[G].$ (2)
Evidently, Eq. (1) defines a one-to-one map between any vector $\boldsymbol{x}\in F^{\ell}$ with coefficients $x_{g}$ labeled by group elements and a group algebra element $x\in F[G]$, and a related map between sets of vectors and sets of group algebra elements. A left $G$-code in $F^{\ell}$ is such a map of a left ideal $J_{L}$ in the ring $F[G]$, defined as an $F$-linear space of elements of $F[G]$ such that for any $x\in J_{L}$ and any $r\in F[G]$, $rx\in J_{L}$. A right $G$-code is defined similarly in terms of a right ideal $J_{R}$, with the opposite order in the product, $xr\in J_{R}$ for any $x\in J_{R}$ and any $r\in F[G]$.
The structure of ideals in $F[G]$ is particularly simple if characteristic of the field does not divide the group size, $\mathop{\rm gcd}(p,\ell)=1$. Then, according to Maschke’s theorem, the group algebra is semisimple, and any ideal is a principal ideal generated by an idempotent, e.g., $J_{L}=F[G]\cdot f_{J}$ for a left ideal, with idempotent $f_{J}^{2}=f_{J}\in J_{L}$, and similarly, $J_{R}=e_{J}\cdot F[G]$ for a right ideal, with idempotent $e_{J}^{2}=e_{J}\in J_{R}$ (see, e.g., Corollary 2.2.5 in Ref. 36).
The usual inner product in $F^{\ell}$ is related to the group trace with the help of a linear map[37] ^$\,\boldsymbol{\widehat{\ }}:F[G]\to F[G]$,
$\widehat{a}\equiv\sum_{g\in G}a_{g^{-1}}g=\sum_{g\in G}a_{g}g^{-1}.$ (3)
Namely, for any $\boldsymbol{a},\boldsymbol{b}\in F^{\ell}$, and the corresponding group algebra elements $a,b\in F[G]$,
${\boldsymbol{a}}\cdot{\boldsymbol{b}}\equiv\sum_{g\in G}a_{g}b_{g}=\mathop{\rm Tr}\nolimits_{G}(\widehat{a}b)=\mathop{\rm Tr}\nolimits_{G}(b\widehat{a}).$ (4)
As a reminder, the group trace is defined as the coefficient of the group identity element $1\in G$: for any $a\in F[G]$, $\mathop{\rm Tr}\nolimits_{G}(a)\equiv a_{1}\in F$.
Given a right group code in a semisimple group algebra $F[G]$ equivalent to a right ideal $J_{R}\equiv\widehat{a}\cdot F[G]$ generated by an element $\widehat{a}\in F[G]$, any group algebra element $x$ corresponding to a vector $\boldsymbol{x}$ in the orthogonal code satisfies[37] the equation $xa=0$. If we denote an idempotent $e_{a}^{2}=e_{a}\in F[G]$ such that $e_{a}a=a$, the solution of the orthogonality equation is the left ideal $J_{L}\equiv F[G]\cdot(1-e_{a})$. More generally, a right ideal $J_{R}\leq F[G]$ is called checkable if there is an element $a\in F[G]$ such that[37] $J_{R}=\{x|x\in F[G],a\cdot x=0\}$.
A chain complex is a sequence of abelian groups and homomorphisms between pairs of consecutive groups such that the image of each homomorphism be included in the kernel of the next. Here we will be concerned with the special case of chain complexes of finite-dimensional vector spaces $\ldots,\mathcal{A}_{j-1},\mathcal{A}_{j},\ldots$ over a finite field $F=\mathbb{F}_{q}$, where $q=p^{m}$ is a power of a prime $p$, and we denote the space dimensions $n_{j}\equiv n_{j}({\cal A})=\dim{\cal A}_{j}$. In this case the boundary operators are linear transformations $\partial_{j}:{\cal A}_{j-1}\leftarrow{\cal A}_{j}$ that map between each pair of neighboring spaces, with the requirement $\partial_{j}\partial_{j+1}=0$, $j\in\mathbb{Z}$. We define a $D$-complex ${\cal A}\equiv{\cal K}(A_{1},\ldots,A_{D})$, a bounded chain complex which only contains $D+1$ non-trivial spaces with fixed bases, in terms of $n_{j-1}\times n_{j}$ matrices $A_{j}$ over $F$ serving as the boundary operators, $j\in\{1,\ldots,D\}$:
${\cal A}:\;\ldots\leftarrow\{0\}\stackrel{{\scriptstyle\partial_{0}}}{{\leftarrow}}{\cal A}_{0}\stackrel{{\scriptstyle A_{1}}}{{\leftarrow}}{\cal A}_{1}\ldots\stackrel{{\scriptstyle A_{D}}}{{\leftarrow}}{\cal A}_{D}\stackrel{{\scriptstyle\partial_{D+1}}}{{\leftarrow}}\{0\}\ldots$ (5)
Here the products of neighboring matrices must be zero, $A_{j-1}A_{j}=0$, $j\in\{2,\ldots,D\}$. In addition to boundary operators given by the matrices $A_{j}$, implicit are the trivial operators $\partial_{0}:\{0\}\leftarrow{\cal A}_{0}$ and $\partial_{D+1}:{\cal A}_{D}\leftarrow\{0\}$ (with the image being the zero vector in ${\cal A}_{D}$) treated formally as rank-zero $0\times n_{0}$ and $n_{D}\times 0$ matrices.
Elements of the subspace $\mathop{\rm im}(\partial_{j+1})\subseteq{\cal A}_{j}$ are called boundaries; in our case these are linear combinations of columns of $A_{j+1}$ and, therefore, form an $F$-linear code with the generator matrix $A_{j+1}^{T}$, $\mathop{\rm im}(A_{j+1})=\mathcal{C}_{A_{j+1}^{T}}$. In the singular case $j=D$, $\mathop{\rm im}(\partial_{D+1})=\{0\}$, a trivial vector space. Elements of $\mathop{\rm Ker}(\partial_{j})\subset{\cal A}_{j}$ are called cycles; in our case these are vectors in an $F$-linear code with the parity check matrix $A_{j}$, $\mathop{\rm Ker}(A_{j})=\mathcal{C}_{A_{j}}^{\perp}$. In the singular case $j=0$, $\mathop{\rm Ker}(\partial_{0})={\cal A}_{0}$, the entire space.
Because of the orthogonality $\partial_{j}\partial_{j+1}=0$, all boundaries are necessarily cycles, $\mathop{\rm im}(\partial_{j+1})\subseteq\mathop{\rm Ker}(\partial_{j})\subseteq{\cal A}_{j}$. The structure of the cycles in ${\cal A}_{j}$ that are not boundaries is described by the $j$ th homology group,
${H}_{j}({\cal A})\equiv H(A_{j},A_{j+1})=\mathop{\rm Ker}(A_{j})/\mathop{\rm im}(A_{j+1}).$ (6)
Group quotient here means that two cycles [elements of $\mathop{\rm Ker}(A_{j})$] that differ by a boundary [element of $\mathop{\rm im}(A_{j+1})$] are considered equivalent; non-zero elements of $\mathcal{H}_{j}(\mathcal{A})$ are equivalence classes of homologically non-trivial cycles. Explicitly, the equivalence of $\boldsymbol{x}$ and $\boldsymbol{y}$ in $\mathcal{A}_{j}$ implies that for some $\boldsymbol{\alpha}\in{\cal A}_{j+1}$, $\boldsymbol{y}=\boldsymbol{x}+\boldsymbol{\alpha}A_{j+1}^{T}$. The rank of $j$-th homology group is the dimension of the corresponding vector space; one has
$k_{j}\equiv\mathop{\rm rank}H_{j}(\mathcal{A})=n_{j}-\mathop{\rm rank}A_{j}-\mathop{\rm rank}A_{j+1}.$ (7)
The homological distance $d_{j}$ is the minimum Hamming weight of a non-trivial element (any representative) in the homology group $H_{j}(\mathcal{A})\equiv H(A_{j},A_{j+1})$,
$d_{j}=\min_{0\not\simeq x\in H_{j}(\mathcal{A})}\mathop{\rm wgt}\nolimits x=\min_{x\in\mathop{\rm Ker}({A}_{j})\setminus\mathop{\rm im}(A_{j+1})}\mathop{\rm wgt}\nolimits x.$ (8)
By this definition, $d_{j}\geq 1$. To address singular cases, throughout this work we define the minimum of an empty set as infinity; $k_{j}=0$ is always equivalent to $d_{j}=\infty$. In particular, the distance of the homology group $H_{0}(\mathcal{A})$ is $d_{0}=1$, unless $A_{1}$ has full row rank, giving $k_{0}=0$, in which case we get $d_{0}=\infty$. In the case of the homology group $H_{D}(\mathcal{A})$, the distance $d_{D}$ is that of the $F$-linear code with the check matrix $A_{D}$.
In addition to the homology group $H(A_{j},A_{j+1})$, there is also a co-homology group $\tilde{H}_{j}(\tilde{\cal A})=H(A_{j+1}^{T},A_{j}^{T})$ of the same rank (7); this is associated with the co-chain complex $\tilde{\cal A}$ formed from the transposed matrices $A_{j}^{T}$ taken in the opposite order. A quantum CSS code[38, 39] with generator matrices $G_{X}=A_{j}$ and $G_{Z}=A_{j+1}^{T}$ is isomorphic with the direct sum of the groups $H_{j}$ and $\tilde{H}_{j}$,
$\mathop{\rm CSS}\nolimits(A_{j},A_{j+1}^{T})\cong H(A_{j},A_{j+1})\oplus H(A_{j+1}^{T},A_{j}^{T}).$ (9)
The two terms correspond to $Z$ and $X$ logical operators, respectively, with stabilizer generator matrices $H_{X}=A_{j}$ and $H_{Z}=A_{j+1}^{T}$. This gives for the homological distances in the chain complex and in the co-chain complex, respectively, $d_{j}=d_{Z}$ and $\tilde{d}_{j}=d_{X}$.
Quantum codes operate in a fault-tolerant (FT) regime, where errors may happen during the syndrome measurement. Commonly used in such a situation is the tactic of redundant measurements. Assuming each operator being measured has the corresponding row in $H_{X}$ or $H_{Z}$, redundancy implies the existence of non-trivial matrices $M_{X}$ and $M_{Z}$ such that $M_{X}H_{X}=0$, $M_{Z}H_{Z}=0$. Rows of $M_{X}$ and $M_{Z}$, respectively, correspond to $X$ and $Z$ metachecks[14]. Evidently, a quantum CSS code with redundant measurements forms a $4$-chain complex with boundary operators $M_{X}$, $H_{X}$, $H_{Z}^{T}$, $M_{Z}^{T}$. In particular, Campbell et al.[14, 22] argued that such longer complexes, e.g., constructed as tensor products of $1$-complexes, have nice error-correction properties in the FT regime, in particular, they can be used for single-shot error correction.
Tensor product $\mathcal{A}\times\mathcal{B}$ of two chain complexes $\mathcal{A}$ and $\mathcal{B}$ is defined as the chain complex formed by linear spaces decomposed as direct sums of Kronecker products,
$(\mathcal{A}\times\mathcal{B})_{j}=\bigoplus\nolimits_{i\in\mathbb{Z}}\mathcal{A}_{i}\otimes\mathcal{B}_{j-i},$ (10)
with the action of the boundary operators
′′′$\partial^{\prime\prime\prime}(a\otimes b)\equiv\partial^{\prime}a\otimes b+(-1)^{i}a\otimes\partial^{\prime\prime}b,$ (11)
where $a\in\mathcal{A}_{i}$, $b\in\mathcal{B}_{j-i}$, and the boundary operators $\partial^{\prime}$, $\partial^{\prime\prime}$, and $\partial^{\prime\prime\prime}$ act in complexes $\mathcal{A}$, $\mathcal{B}$, and $\mathcal{A}\times\mathcal{B}$, respectively. Notice that the two terms in Eq. (11) are supported in different subspaces of the expansion (10). When both $\mathcal{A}$ and $\mathcal{B}$ are bounded, the dimension $n_{j}(\mathcal{C})$ of a space $\mathcal{C}_{j}$ in the product $\mathcal{C}=\mathcal{A}\times\mathcal{B}$ is
$n_{j}(\mathcal{C})=\sum\nolimits_{i}n_{i}(\mathcal{A})\,n_{j-i}(\mathcal{B}).$ (12)
The homology groups of the product $\mathcal{C}={\cal A}\times{\cal B}$ are isomorphic to a simple expansion in terms of those of ${\cal A}$ and ${\cal B}$ which is given by the Künneth formula,
$H_{j}(\mathcal{C})\cong\bigoplus\nolimits_{i}H_{i}(\mathcal{A})\,\otimes\,H_{j-i}(\mathcal{B}).$ (13)
One immediate consequence is that the rank $k_{j}(\mathcal{C})$ of the $j$ th homology group $H_{j}(\mathcal{C})$ is
$k_{j}(\mathcal{C})=\sum\nolimits_{i}k_{i}(\mathcal{A})\,k_{j-i}(\mathcal{B}).$ (14)
Such a convolution can be also written as a product of the Poincare polynomials $p_{\mathcal{A}}(x)\equiv\sum_{j}k_{j}(\mathcal{A})x^{j}$ corresponding to the two complexes, $p_{\mathcal{C}}(x)=p_{\mathcal{A}}(x)p_{\mathcal{B}}(x)$.
For the homological distances in a product of two arbitrary-length complexes, we have some bounds[21],
$\min_{i\in\mathbb{Z}}\max\biglb(d_{i}(\mathcal{A}),d_{j-i}(\mathcal{B})\bigrb)\leq d_{j}(\mathcal{C})\leq\min_{i\in\mathbb{Z}}d_{i}(\mathcal{A})d_{j-i}(\mathcal{B}).$ (15)
The exact distances are known only in the important case where one of the factors is a $1$-complex[35, 20, 21]. With a $D$-complex $\mathcal{A}$ in Eq. (5) and a $1$-complex $\mathcal{B}\equiv\mathcal{K}(B_{1})$, the homological distances are[21]
$d_{j}(\mathcal{C})=\min\biglb(d_{j}(\mathcal{A})d_{0}(\mathcal{B}),d_{j-1}(\mathcal{A})d_{1}(\mathcal{B})\bigrb).$ (16)
In the following we refer to a $D$-fold product of $1$-complexes as a $D$-dimensional QHP complex[35, 20, 21]. Such a $D$-complex ${\cal Q}$ has $D+1$ non-trivial linear spaces ${\cal Q}_{j}$, with $0\leq j\leq D$, where the space at level $j$ is formed as a direct sum of $m_{j}^{(D)}$ product subspaces,
${\cal A}^{(1)}_{j_{1}}\times{\cal A}^{(2)}_{j_{2}}\times\ldots\times{\cal A}^{(D)}_{j_{D}}.$
Here the level index $j=j_{1}+j_{2}+\ldots+j_{D}$ is a sum of individual level indices $j_{i}\in\{0,1\}$; evidently, the number of such terms is given by a binomial,
$m_{j}^{(D)}={D\choose j}.$ (17)
In particular, consider a $1$-complex ${\cal A}\equiv\mathcal{K}(A)$, where $A$ is a square $n\times n$ matrix such that an $F$-linear code ${\cal C}_{A_{1}}^{\perp}$ has parameters $[n,k,d]$. Its $D$-fold tensor product with itself, $\mathcal{Q}={\cal A}^{\times D}$, has space dimensions
$n_{j}^{(D)}={D\choose j}n^{D},\quad 0\leq j\leq D,$ (18)
and the dimension and the homological distance of its $j$ th homology group $H_{j}(\mathcal{Q})$ are, respectively,
$k_{j}^{(D)}={D\choose j}k^{D},\quad d_{j}^{(D)}=d^{j},\quad 0\leq j\leq D.$ (19)
III Construction and code properties.
III.1 Multi-block complexes and associated quantum codes
We construct a multi-block chain (MBC) complex as a generalization of $D$-dimensional QHP codes in exactly the same fashion as the two-block codes have been constructed as a generalization of regular (2D) HP codes[24]. Namely, given matrices $A_{1}$ and $B_{1}$ over $F$, with dimensions $n_{0}(\mathcal{A})\times n_{1}(\mathcal{A})$ and $n_{0}(\mathcal{B})\times n_{1}(\mathcal{B})$, respectively, the tensor product $\mathcal{P}=\mathcal{A}\times\mathcal{B}$ of the complexes $\mathcal{A}\equiv\mathcal{K}(A_{1})$ and $\mathcal{B}\equiv\mathcal{K}(B_{1})$, explicitly,
$\mathcal{P}:\ldots\{0\}\stackrel{{\scriptstyle\partial_{0}}}{{\leftarrow}}{\cal A}_{0}\times{\cal B}_{0}\stackrel{{\scriptstyle P_{1}}}{{\leftarrow}}\begin{array}[c]{c}{\cal A}_{0}\times{\cal B}_{1}\\ {\cal A}_{1}\times{\cal B}_{0}\end{array}\stackrel{{\scriptstyle P_{2}}}{{\leftarrow}}{\cal A}_{1}\times{\cal B}_{1}\leftarrow\{0\}\ldots,$
has the boundary operators [cf. Eq. (11)]
${P}_{1}=\left[I_{A}\otimes B_{1},A_{1}\otimes I_{B}\right],\quad{P}_{2}=\left[\begin{array}[c]{c}A_{1}\otimes I_{B}\\ -I_{A}\otimes B_{1}\end{array}\right],$ (20)
where $I_{A}\equiv I(\mathcal{A}_{0})$ and $I_{B}\equiv I(\mathcal{B}_{0})$ are the identity operators on the spaces $\mathcal{A}_{0}$ and $\mathcal{B}_{0}$, respectively. The product vanishes, $P_{1}P_{2}=0$, because the two blocks, $A\equiv A_{1}\otimes I_{B}$ and $B\equiv I_{A}\otimes B_{1}$, commute. The generalization[24] is evident: for any pair of $\ell\times\ell$ square commuting matrices, $AB=BA$, the two-block code has the CSS matrices
$H_{X}\equiv P_{1}=\left[B,A\right],\quad H_{Z}^{T}\equiv P_{2}=\left[\begin{array}[c]{c}A\\ -B\end{array}\right].$ (21)
We regard this construction as the smallest (two-dimensional) multi-block chain complex, and denote it ${\cal P}=\mathop{\text{\small\sc MBC}}\nolimits(A,B)$, with the requirement $AB=BA$.
Similarly, to define a three-dimensional generalization, a 3-chain complex $\mathop{\text{\small\sc MBC}}\nolimits(A,B,C)$ from three pairwise commuting matrices, we insert an additional $1$-complex into the product, $\mathcal{Q}\equiv\mathcal{A}\times\mathcal{B}\times\mathcal{C}=\mathcal{P}\times{\cal C}$, with the non-trivial subspaces
$\mathcal{P}_{0}\times\mathcal{C}_{0}$ $\stackrel{{\scriptstyle Q_{1}}}{{\leftarrow}}$ $\begin{array}[c]{c}\mathcal{P}_{0}\times\mathcal{C}_{1}\\ \mathcal{P}_{1}\times\mathcal{C}_{0}\\ \end{array}\stackrel{{\scriptstyle Q_{2}}}{{\leftarrow}}\begin{array}[c]{c}\mathcal{P}_{1}\times\mathcal{C}_{1}\\ \mathcal{P}_{2}\times\mathcal{C}_{0}\end{array}\stackrel{{\scriptstyle Q_{3}}}{{\leftarrow}}\mathcal{P}_{2}\times\mathcal{C}_{1}.$
The subspaces in direct sums are stacked vertically [cf. Eq. (10)], and ordered alphabetically by the space indices. The structure of the corresponding boundary operators follows directly from the subspace decomposition and the sign convention (11), $Q_{1}=\left[I\otimes C_{1},P_{1}\otimes I\right]$,
$Q_{2}=\left[\begin{array}[c]{cc}P_{1}\otimes I&0\\ -I\otimes C_{1}&{P}_{2}\otimes I\end{array}\right],\;\,\text{and}\,\;Q_{3}=\left[\begin{array}[c]{c}P_{2}\otimes I\\ I\otimes C_{1}\end{array}\right].$
With Eq. (20), it is easy to verify that the resulting matrices contain only three blocks, $A\equiv A_{1}\otimes I_{B}\otimes C_{1}$, $B\equiv I_{A}\otimes B_{1}\otimes I_{C}$, and $C\equiv I_{A}\otimes I_{B}\otimes C_{1}$, and can be explicitly written as
$Q_{1}$ $=$ $[C\,|\,B\;A],$ (23)
$Q_{2}$ $=$ $\left[\begin{array}[c]{cc|c}B&A&0\\ \hline\cr-C&0&A\\ 0&-C&-B\end{array}\right],$ (27)
$Q_{3}$ $=$ $\left[\begin{array}[c]{c}A\\ -B\\ \hline\cr C\end{array}\right].$ (31)
We used the lines here to emphasize the inherited block structure, cf. matrices in Eq. (21). Just like for the case of two-block codes, the orthogonality can be verified directly, $Q_{1}Q_{2}=0$, $Q_{2}Q_{3}=0$, as long as the individual blocks $A$, $B$, and $C$ commute.
For a general definition, consider a $D$-complex ${\cal Q}\equiv\mathop{\text{\small\sc MBC}}\nolimits(A_{1},A_{2},\ldots,A_{D})$ constructed from $D$ commuting square blocks $A_{i}$, $\ell\times\ell$ matrices over $F$, with boundary operators $\partial_{j}^{(Q)}=Q_{j}$, where $Q_{j}Q_{j+1}=0$. Explicitly, the space dimensions (non-zero for $0\leq j\leq D$) are $\dim({\cal Q}_{j})\equiv n_{j}^{(Q)}={D\choose j}\ell$; it is also convenient to define the block dimensions $m_{j}\equiv m_{j}^{(Q)}={D\choose j}$ such that $n_{j}=m_{j}\ell$. Then, given an additional block $A_{D+1}\equiv N$ commuting with those already present, the $(D+1)$-dimensional chain complex ${\cal R}\equiv\mathop{\text{\small\sc MBC}}\nolimits(A_{1},A_{2},\ldots,A_{D},N)$ has the block dimensions $m_{j}^{(R)}=m_{j-1}^{(Q)}+m_{j}^{(Q)}={D+1\choose j}$, and the boundary operators constructed as block matrices,
$R_{1}$ $=$ $\bigl{[}\begin{array}[]{cc}N,&Q_{1}\end{array}\bigr{]},$ (33)
[…]
$R_{D+1}$ $=$ $\left[\begin{array}[]{c}Q_{D}\\ (-1)^{D}N\end{array}\right],$ (39)
where $I(m)$ is an $m\times m$ identity matrix.
For a $D$-dimensional MBC complex with boundary matrices $Q_{j}$, $1\leq j\leq D$, it is easy to verify that all $D$ blocks are used in an equivalent fashion, i.e., all chain complexes constructed from the same $D$ commuting blocks taken in different orders are equivalent (can be obtained from each other by block row and column permutations). In particular, matrix $Q_{j}$ has $j$ non-zero blocks in each column and $D+1-j$ non-zero blocks in each row. Furthermore, in each pair of matrices $Q_{j}$ and $Q_{j+1}^{T}$ with $m_{j}={D\choose j}$ column blocks, $1\leq j<D$, each pair of matching column-blocks contains a full set of blocks $\{A_{i}\}$, exactly $j$ such blocks in $Q_{j}$ and the remaining $D-j$ (transposed) blocks in $Q_{j+1}^{T}$.
For future reference, the boundary operators in the 4-complex ${\cal R}\equiv\mathop{\text{\small\sc MBC}}\nolimits(A,B,C,D)$ are
$R_{1}$ $=$ $\,\biglb[D\,|\,C\,B\,A\bigrb]\,,$ (40)
$R_{2}$ $=$ $\left[\begin{array}[c]{ccc|ccc}C&B&A\\ \hline\cr-D&&&B&A&0\\ &-D&&-C&0&A\\ &&-D&0&-C&-B\\ \end{array}\right],$ (45)
$R_{3}$ $=$ $\left[\begin{array}[c]{ccc|c}B&A&0\\ -C&0&A\\ 0&-C&-B\\ \hline\cr D&&&A\\ &D&&-B\\ &&D&C\\ \end{array}\right],\;\,\,$ (52)
$R_{4}$ $=$ $\left[\begin{array}[c]{ccc|c}A\\ -B\\ C\\ \hline\cr-D\end{array}\right].$ (57)
While any non-trivial homology group in this complex may be used to construct a quantum code, we will focus on ${H}_{2}(\mathcal{R})=H(R_{2},R_{3})$ and the corresponding co-homology group $\tilde{H}_{2}({\cal R})$. This gives a self-dual CSS code with generator matrices $H_{X}=R_{2}$, $H_{Z}=R_{3}^{T}$.
III.2 Abelian multi-cycle (AMC) codes.
AMC complexes (and associated quasi-abelian quantum CSS codes) are a special case of multi-block codes where the commuting matrices are constructed [^1] Commuting matrices can be constructed for any pair of group algebra elements even for a non-abelian group[58]. While there is no direct generalization for arbitrary sets of $D>2$ non-abelian group algebra elements, less general constructions do exist. These go outside the scope of the present work. with the help of an abelian group algebra $F[G]$, see Eq. (1). Namely, assuming the group size is $\ell\equiv|G|$, for an element $a\in F[G]$, the $\ell\times\ell$ matrix $A\equiv\mathop{\mathrm{M}}\nolimits(a)$ is defined by the action on group elements,
$[\mathop{\mathrm{M}}\nolimits(a)]_{\alpha,\beta}\equiv\sum_{g\in G}a_{g}\delta_{\alpha,g\beta},$ (58)
where group elements $\alpha,\beta\in G$ are used to index rows and columns, and $\delta_{\alpha,\beta}=1$ if $\alpha=\beta$ and 00 otherwise is the Kronecker delta. It is easy to verify that for group elements $g\in G$, matrices $\mathop{\mathrm{M}}\nolimits(g)$ form the regular $F$-linear representation of $G$.
Then, given a sequence $\{a_{j}\}_{j=1}^{D}\subset F[G]$ of abelian group algebra elements, with $D\geq 2$, the corresponding matrices $A_{j}=\mathop{\mathrm{M}}\nolimits(a_{j})$ commute, and the $D$-complex $\mathcal{Q}\equiv\mathcal{Q}^{(D)}=\mathop{\text{\small\sc MBC}}\nolimits(A_{1},A_{2},\ldots,A_{D})$ can be constructed as explained in the previous section. We denote such a $D$-complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$. The corresponding level-$j$ quantum CSS code $\mathop{\text{\small\sc AMC}}\nolimits_{j}(a_{1},a_{2},\ldots,a_{D})$, $j\leq D$, is constructed from the level-$j$ homology $H_{j}(\mathcal{Q})$ and co-homology $H_{j}(\mathcal{Q})$ groups. Explicitly,
$\mathop{\text{\small\sc AMC}}\nolimits_{j}(a_{1},a_{2},\ldots,a_{D})=\mathop{\rm CSS}\nolimits(Q_{j-1},Q_{j}^{T}),$
where $Q_{j}$ is the $j$-th boundary operator (matrix) in the $D$-complex $\mathcal{Q}$.
To visualize, consider a finite abelian group
$G=\langle x_{1},x_{2},\ldots,x_{m}|r_{1}(\boldsymbol{x}),r_{2}(\boldsymbol{x}),\ldots,r_{m}(\boldsymbol{x})\rangle_{\rm abelian},$
presented with $m$ commuting generators $x_{1}$, $x_{2}$, …, $x_{m}$, and $m$ additional independent (monomial) relators
$1_{G}=r_{i}(\boldsymbol{x})=\prod_{j=1}^{m}x_{j}^{\Delta_{j}},\quad\Delta_{j}\equiv{\Delta_{j}^{(i)}},\quad i=1,2,\ldots,m.$
Further, elements of the free abelian group $\langle x_{1},x_{2},\ldots,x_{m}\rangle$ are in a one-to-one correspondence with points of $\mathbb{Z}^{m}$. The exponents in each relator $r_{i}(\boldsymbol{x})$ form a basis vector $\boldsymbol{\Delta}^{(i)}=(\Delta_{1}^{(i)},\Delta_{2}^{(i)},\ldots,\Delta_{m}^{(i)})$ of a lattice on $\mathbb{Z}^{m}$; individual basis vectors are combined as columns of a lattice basis matrix $\Delta$. Any two points on $\mathbb{Z}^{m}$ connected by a lattice vector are identified. Then, elements of the group $G$ are in a one-to-one correspondence with inequivalent points in $\mathbb{Z}^{m}$ which form an $m$-torus ${\cal T}_{m}$. The number of inequivalent points (also, the number of points in a primitive cell of the lattice) is $n=\left|G\right|=\left|\det\Delta\right|$. In other words, $\mathcal{T}_{m}$ is covered by $\mathbb{Z}^{m}$, with a well-defined covering function $f:\mathbb{Z}^{m}\to{\cal T}_{m}$ which is one-to-one in each ball of some radius $R>0$ on $\mathbb{Z}^{m}$, the injectivity radius.
Respectively, support of an element $a\equiv a(\boldsymbol{x})\in F[\boldsymbol{x}]$, an $m$-variate polynomial over $F$, corresponds to a set of points on ${\cal T}_{m}$; multiplication by a generator $x_{j}$ gives a translation along the $j$-axis. A presentation where each relator depends on one variable only, $r_{i}=x_{i}^{\Delta_{i}}$, corresponds to a decomposition [^2] Such a decomposition always exists according to the Fundamental Theorem of finite abelian groups. of a finite abelian group as a direct product of cyclic groups,
$G=C_{\Delta_{1}}\times C_{\Delta_{2}}\times\ldots\times C_{\Delta_{m}}.$
In such a case we get a torus with periodicity vectors along the Cartesian axes, and the group order is given by a product, $|G|=\prod_{i=1}^{m}\Delta_{i}$.
We note that the number of generators in a presentation of the same group may vary. In particular, while any cyclic group $C_{n}=\langle x|x^{n}\rangle$ has a single-generator presentation, presentations with two or more variables can always be constructed for a sufficiently large group. For example,
$C_{15}$ $=$ $C_{3}\times C_{5}=\langle x,y|x^{3},y^{5}\rangle_{\rm abelian},\quad\Delta={\small\left(\begin{array}[c]{cc}3,&0\\ 0,&5\end{array}\right)},$
$C_{15}$ $=$ $\langle x,y|x^{4}y,xy^{4}\rangle_{\rm abelian},\quad\Delta={\small\left(\begin{array}[c]{cc}4,&1\\ 1,&4\end{array}\right).}$
where the only condition is that the determinant of the matrix $\Delta$ matches the group order, $n=\left|\det\Delta\right|$.
III.3 Equivalent AMC codes
Changing a group presentation, one can try to come up with an equivalent complex, e.g., with better locality properties. How much freedom do we have? To answer this question, let us first discuss the symmetries of AMC construction. The following statement is given without a proof, as a generalization of Theorem 3 in Ref. 28:
Statement 1.
For any sequence $\{a_{j}\}\equiv\{a_{j}\}_{j=1}^{D}\subset F[G]$, the $D$-complex $\mathop{\text{\small\sc AMC}}\nolimits(\{a_{j}\})$ is permutation-equivalent to the $D$-complex constructed from the modified sequence:
  1. (i)
    A permuted sequence of group algebra elements;
  2. (ii)
    $\{\varphi(a_{j})\}$, for any automorphism $\varphi:G\to G$;
  3. (iii)
    $\{x_{j}a_{j}\}$, for any non-zero $x_{j}\in F$;
  4. (iv)
    $\{a_{j}\alpha_{j}\}$, for any $\alpha_{j}\in G$.
In addition, the sequence $\{\widehat{a}_{j}\}$, see Eq. (3), gives an AMC complex isomorphic to the original co-chain complex, $\mathop{\text{\small\sc AMC}}\nolimits(\{\widehat{a}_{j}\})\simeq\widetilde{\mathop{\text{\small\sc AMC}}\nolimits}(\{a_{j}\})$.
Therefore, we can always “rescale” the elements in the original set, and choose $a_{j}=1+\ldots$, $j\leq D$. Further, we can try to construct a $D$-local complex in $\mathbb{Z}^{D}$, e.g., by choosing group generators in the support of each element, e.g., $a_{j}=1+x_{j}+(\text{other monomials})$. Clearly, the second transformation changes the group representation and the basis vectors associated with the lattice.
For example, with weight-two elements $a_{j}=1+x_{j}$, $1\leq j\leq D$, we get a complex locally equivalent to $D$-dimensional QHP codes. Assuming the entire group can be generated this way [^3] Otherwise, the spaces in an AMC complex can be decomposed onto a direct sum of subspaces corresponding to cosets of the subgroup $G_{X}=\langle\{x_{j}\}\rangle$ in the original group $G$; see Sec. IV-C in Ref. 28. In the abelian case these complexes are permutation-equivalent to a complex in the subgroup $G_{X}\leq G$.,
$G=\left\langle\{x_{j}\}_{j=1}^{D}\right\rangle,$
we get a $D$-complex locally equivalent to $D$-dimensional toric codes. With a level-$j$ quantum CSS code, for errors with irreducible weight smaller than the injectivity radius $R$, this guarantees identical syndrome weights, that is, identical confinement profiles[22] up to $R$. We also expect the code distance to grow as $d={\cal O}(R^{\alpha})$, with the exponent $\alpha=\min(j,D-j)$ expected for level-$j$ $D$-dimensional toric codes, see Statement 4 below. Further, we can use a measuring circuit designed for a regular level-$j$ $D$-dimensional toric code: its validity is guaranteed as a local property. [^4] Note, however, that the circuit distance need not always coincide with that of the original code. As an example, while any single-ancilla measurement circuit would work for a toric code[59], rotated surface codes require a carefully designed $\sf Z$-$\sf N$ addressing scheme for fault-tolerance[51].
In practice, especially if we want to search for AMC chain complexes constructed from group algebra elements with weights higher than two and groups with two or more generators, it may be more efficient to go over small-weight polynomials and suitable periodicity vectors. For example, such a method was used in Ref. 30 to construct a number of weight-six locally-planar GB (also known as BB) codes.
III.4 Matrix ranks for AMC construction.
While the ranks of the matrices are easy to compute numerically in each particular case (e.g., with the help of a computer algebra system capable of doing linear algebra over the field $F$), we have analytically computed the corresponding ranks in two important cases.
First, consider a cyclic group of order $\ell$, with the group algebra $F[C_{\ell}]$ isomorphic to the ring $F[x]/(x^{\ell}-1)$ of modular polynomials. Here the ranks of the boundary operators in the $D$-complex $\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$ constructed from single-variate polynomials $a_{j}\equiv a_{j}(x)$, $1\leq j\leq D$, is expressed in terms of the characteristic polynomial,
$h(x)=\mathop{\rm gcd}(a_{1},a_{2},\ldots,a_{D},x^{\ell}-1).$ (61)
In Appendix A, we prove the following statement:
Statement 2.
Consider a $D$-complex $\mathop{\text{\small\sc AMC}}\nolimits(\{a_{j}\}_{j=1}^{D})$ over $F[C_{\ell}]$. The boundary operators have $\mathop{\rm rank}Q_{j}={D-1\choose j-1}(\ell-\kappa)$, where $\kappa=\deg h(x)$ is the degree of the characteristic polynomial (61).
As an immediate consequence, the corresponding homology groups have orders $\left|H_{j}({\cal Q})\right|={D\choose j}\kappa$. Previously, the same or equivalent results were derived for the special case of $D=2$ AMC codes (quasi-cyclic GB codes), see Refs. [24, 44]. Also note a similarity with the corresponding equations for $D$-dimensional HGP codes. The main difference with Eqs. (18) and (19) is the linear scaling with the group order $\ell$, which is why the AMC construction gives much smaller codes.
Second is the case of a semisimple group algebra $F[G]$, where the group order $\ell=\left|G\right|$ is mutually prime with the field characteristic $p$, $\mathop{\rm gcd}(p,\ell)=1$. In this case the result is very similar to that in Statement 2, except the value of $\kappa$ is defined differently. For a technical definition of $\kappa$ in terms of products of idempotent matrices, please see the proof given in Appendix B. In practice, $\kappa$ can be defined, e.g., as the dimension of the classical $F$-linear code with the check matrix ${Q}_{D}$ (all $D$ square blocks stacked together in a single column block).
Statement 3.
Consider a $D$-dimensional complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$ over a semi-simple abelian group algebra $F[G]$. The ranks of all boundary operators are proportional to that of $Q_{D}$, $\mathop{\rm rank}Q_{j}={D-1\choose j-1}\mathop{\rm rank}Q_{D}$.
The same result can be also proven using a decomposition of the semisimple abelian group $G$ into a product of semisimple cyclic groups, see the Appendix in Ref. 45.
III.5 Distance bounds.
First we construct lower (existence) bounds on homological distances in AMC complexes. They are based on the results for $D$-dimensional QHP codes[35, 20, 21], and the fact that the AMC construction includes QHP codes derived from abelian group-algebra codes as a special case. We prove in Appendix C:
Statement 4.
Let $G$ be a finite abelian group presented with the list of generators $\boldsymbol{x}=(x_{1},x_{2},\ldots,x_{m})$, and $K\equiv G^{\times D}$ its $D$-fold tensor product with itself, with the corresponding list of generators $\{\boldsymbol{x}_{1},\boldsymbol{x}_{2},\ldots,\boldsymbol{x}_{D}\}$. Given a checkable[37] abelian group-algebra code in $F[G]$ of length $n$, dimension $k$, and distance $d$, with the orthogonal element $a\equiv a(\boldsymbol{x})$, consider the group algebra elements $a_{i}\equiv a(\boldsymbol{x}_{i})\in F[K]$, $1\leq i\leq D$, and the corresponding $D$-complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(a_{1},a_{2},\ldots,a_{D})$. Its $j$-th subspace has dimension $\dim({\cal Q}_{j})={D\choose j}n^{D}$, while the corresponding homology and co-homology groups have the same dimension $k_{j}={D\choose j}k^{D}$ and homological distances $d_{j}=d^{j}$ and $\tilde{d}_{j}=d^{D-j}$, respectively.
In the special case of symmetric even-$D$ AMC codes at level $j=D/2$, this gives the same rate as the original code, and the distance $d^{D/2}$. Thus, a finite rate and a linear distance scaling in the original code family would give a finite rate and the distance scaling as a square root of the code length, similar to GB codes. While we are not aware of a proof that good linear abelian group-algebra codes exist (known results include good code families based on non-abelian dihedral groups $D_{2m}$, random quasi-abelian codes [46], and non-linear cyclic codes[47] that attain the Gilbert-Varshamov bound), already the classical results on the cyclic codes[48, 49] are sufficient to show the existence of AMC codes with finite relative distances and power-law distance scaling, with the exponent arbitrarily close to $1/2$.
Second, we give an upper bound on the distance for any AMC code based on a cyclic group. The bound, formulated in terms of the characteristic polynomial (61) is a generalization of a similar result for GB codes[26].
Statement 5.
Take a $D$-complex ${\cal Q}=\mathop{\text{\small\sc AMC}}\nolimits(\{a_{i}\}_{i=1}^{D})$ over $F[C_{\ell}]$ with characteristic polynomial $h(x)$. Let $d_{h}^{\perp}$ be the minimal distance of the $F$-linear cyclic code with the check polynomial $h(x)$. The homological distances satisfy the inequality $d_{j}({\cal Q})\leq d_{D}({\cal Q})=d_{h}^{\perp}$, $1\leq j<D$.
When boundary operators from a level-$j$ homology group are used to construct a CSS code, this gives a uniform upper bound on the code distances, $d\leq d_{h}^{\perp}$. We also get a bound on the single-shot distance[14] $d_{\rm SS}\leq d_{h}^{\perp}$, a parameter which may be relevant when the code is operated in a fault-tolerant setting.
Notice, however, that the single-shot distance $d_{\rm SS}$ is only relevant if data and syndrome decoding are done separately, and that only if we are using unmodified matrices from the chain complex, e.g., $M_{X}=Q_{j-1}$ and $M_{Z}=Q_{j+2}^{T}$ with $H_{X}=Q_{j}$ and $H_{Z}=Q_{j+1}^{T}$. In practice, the single-shot distance may be increased by adding homologically non-trivial chains (co-chains) as additional rows of $M_{X}$ ($M_{Z}$). The downside is that these vectors may have higher weights, potentially increasing the complexity of the syndrome decoding step.
Another parameter relevant in a fault-tolerant setting is the syndrome distance $d_{\rm S}$, the minimum weight of a non-zero syndrome vector. An obvious upper bound is the minimum column weight in the check matrix. A boundary operator $Q_{j}$ in an AMC $D$-complex ${\cal Q}$ has $j$ or $D+1-j$ non-zero blocks in each block column or row, respectively. If we use group algebra elements of the same weight $w$, this gives uniform column weights $jw$ and row weights $(D+1-j)w$. For a code $\mathop{\rm CSS}\nolimits(Q_{j},Q_{j+1}^{T})$ in the space ${\cal Q}_{j}$ of such a $D$-complex, this gives a simple upper bound on the syndrome distance,
$d_{\rm S}\leq\min(j,D-j)w.$ (62)
We also get uniform weights of the stabilizer generators defined by the rows of $H_{X}=Q_{j}$ and $H_{Z}=Q_{j+1}^{T}$,
$w_{X}=(D+1-j)w.\quad w_{Z}=(j+1)w.$ (63)
In particular, with $D=4$ and $j=2$, this gives $d_{\rm S}\leq 2w$ and $w_{X}=w_{Z}=3w$.
IV Numerical results
IV.1 Family of $4$-dimensional rotated toric codes
We now focus on the family of quasicyclic CSS codes $\mathop{\text{\small\sc AMC}}\nolimits_{2}(a_{1},a_{2},a_{3},a_{4})$ constructed from distinct weight-2 group algebra elements $a_{i}\in F[C_{\ell}]$, $1\leq i\leq 4$. As explained in the previous section, such codes are locally equivalent to $4$-dimensional toric codes with qubits on faces (second homology group). While the topology of the code is the same as for the regular $4D$ toric codes, the lattice periodicity vectors $\boldsymbol{\Delta}_{i}$ need no longer be parallel to the Cartesian axes. Thus, we may refer to these codes as rotated $4D$ toric codes.
Specifically, we constructed such codes for $7\leq\ell\leq 30$ and corresponding block lengths $n=6\ell$ (four distinct polynomials cannot be found for $72\leq\ell<$ 7). Generating polynomials and parameters of smallest codes for each minimum distance $d$ are listed in Tab. 1. All constructed codes have the same dimension $k=6$ and the same characteristic polynomial $h(x)=1+x$. For comparison, in Tab. 1 we also list parameters of two conventional $4D$ toric codes, which require larger block lengths to achieve the same distances.
In addition to conventional code parameters and the defining polynomials, Tab. 1 shows the confinement profile[22], a sequence of minimum syndrome weights corresponding to errors of increasing irreducible weights $w\geq 1$. These are important for single-shot error correction in the presence of syndrome measurement errors. Indeed, for a conventional level-2 4D toric code, the unobservable errors are closed surfaces formed by plaquettes, while such a surface with a boundary is observable, where edges along the surface boundary correspond to flipped syndrome bits. With the lattice of size $L^{4}$, the perimeter is expected to scale as $\mathcal{O}(w^{1/2})$, for error weights $w\leq w_{\rm max}=\mathcal{O}(L^{2})$, with the first five terms in the large-$L$ sequence $4,6,6,8,8$. The confinement profiles of the codes in Tab. 1 (computed using the dist_m4ri program [50]) follow this sequence, as long as the corresponding minimum-perimeter configuration does not contain a pair of edges identified by a lattice vector. This is substantially different from $3+3$ GB codes previously studied in Ref. 25, where for all designed codes the confinement profiles (starting with $3,4,3$) are increasing only for irreducible error weights $w\leq 2$.
$\ell$ $n$ $k$ $d$ $d_{\rm S}$ $a_{1}$ $a_{2}$ $a_{3}$ $a_{4}$ confin
7 42 6 4 4 $1+x$ $1+x^{2}$ $1+x^{3}$ $1+x^{4}$ 4,4,4,6
10 60 6 5 4 $1+x$ $1+x^{2}$ $1+x^{3}$ $1+x^{4}$ 4,6,6,6,4
11 66 6 6 4 $1+x$ $1+x^{2}$ $1+x^{3}$ $1+x^{4}$ 4,6,6,6,4
14 84 6 7 4 $1+x$ $1+x^{2}$ $1+x^{5}$ $1+x^{6}$ 4,6,6,6,4
16 96 6 8 4 $1+x$ $1+x^{3}$ $1+x^{5}$ $1+x^{7}$ 4,6,8,8,4
18 108 6 9 4 $1+x$ $1+x^{3}$ $1+x^{5}$ $1+x^{7}$ 4,6,8,8,4
25 150 6 10 4 $1+x$ $1+x^{4}$ $1+x^{6}$ $1+x^{9}$ 4,6,8,8,4
28 168 6 11 4 $1+x$ $1+x^{3}$ $1+x^{7}$ $1+x^{12}$ 4,6,8,8,4
30 180 6 12 4 $1+x^{2}$ $1+x^{5}$ $1+x^{8}$ $1+x^{9}$ 4,6,8,8,4
$2^{4}$ 96 6 4 4 $1+x_{1}$ $1+x_{2}$ $1+x_{3}$ $1+x_{4}$ 4,4,4
$3^{4}$ 486 6 9 4 $1+x_{1}$ $1+x_{2}$ $1+x_{3}$ $1+x_{4}$ 4,6,6,8,4
Table 1: Smallest $D=4$ AMC2 qubit codes (rotated $4D$ toric codes) with increasing distances $d$, constructed from cyclic groups $C_{\ell}$ with $\ell\leq 30$ (above the line) and two conventional $4D$ toric codes (below the line). Confinement profile [22] (“confin”) column lists the minimum syndrome weights for irreducible errors of weights from $1$ to $5$, except for codes with $\ell=7$ and $\ell=2^{4}$ where the biggest irreducible errors have weights $4$ and $3$, respectively.
IV.2 Circuit design
Stabilizer generator matrices of these codes, $H_{X}=R_{2}$ and $H_{Z}=R_{3}^{T}$ are shown in Eqs. (45) and (52), where the blocks $A$, $B$, $C$, and $D$ should be replaced with the corresponding circulant matrices. With weight-two polynomials, this gives stabilizer generators of uniform weight $w=6$. Each data qubit (associated with a plaquette on a hypercubic lattice) is neighboring with four $X$ checks (edges) and four $Z$ checks (cubes), total of $8$ checks.
We have set out to design quantum codes with increased redundancy of low-weight stabilizer generators. Evidently, constructed matrices have too much redundancy to measure all checks in a circuit of optimal duration. For this reason, in each round of measurement we chose to address only the stabilizer generators corresponding to three chosen row blocks, labeling the corresponding circuits by the index of the dropped row block.
As an example, after removing the first row block from the CSS matrices $H_{X}=R_{2}$ and $H_{Z}=R_{3}^{T}$ in Eqs. (45) and (52), we get (we also dropped the signs assuming the binary field)
$H_{X}^{(1)}$ $=$ $\left[\begin{array}[c]{ccc|ccc}D_{1}&&&B_{4}&A_{5}\\ &D_{2}&&C_{4}&&A_{6}\\ &&D_{3}&&C_{5}&B_{6}\end{array}\right],$ (67)
$H_{Z}^{(1)}$ $=$ $\left[\begin{array}[c]{ccc|ccc}\widehat{B}_{1}&\widehat{C}_{2}&&\widehat{D}_{4}\\ \widehat{A}_{1}&&\widehat{C}_{3}&&\widehat{D}_{5}\\ &\widehat{A}_{2}&\widehat{B}_{3}&&&\widehat{D}_{6}\!\end{array}\right],\;\,$ (71)
where the subscripts next to block names label the columns and the circumflex “hat” accent indicates transposed blocks. Focusing on bigger $3\times 3$ blocks, we have a two-block structure similar to that in lifted-product or GB/2BGA codes, with two and four (”$2+4$”) non-zero entries in each subrow. The circuits designed[25] for $2+4$ 2BGA/GB codes, a generalization of the $\sf Z$-$\sf N$ addressing scheme[51] for rotated surface codes, also work in the present case. Namely, a valid time-optimal circuit is obtained if we address the qubits corresponding to the two terms in the $D$ blocks at steps 1 and 6 (first and last), with the opposite order for $H_{X}$ and $H_{Z}$, while addressing the qubits from the remaining blocks in the intervening time steps. Specifically, we used the following addressing scheme:
$\begin{array}[c]{l|ccc|ccc|}H_{X}&D_{10}&B_{40}&B_{41}&A_{50}&A_{51}&D_{11}\\ &D_{20}&A_{60}&A_{61}&C_{40}&C_{41}&D_{21}\\ &D_{30}&C_{50}&C_{51}&B_{60}&B_{61}&D_{31}\\[1.49994pt] \hline\cr H_{Z}&\widehat{D}_{41}&\widehat{B}_{10}&\widehat{B}_{11}&\widehat{C}_{20}&\widehat{C}_{21}&\widehat{D}_{40}\\ &\widehat{D}_{51}&\widehat{C}_{30}&\widehat{C}_{31}&\widehat{A}_{10}&\widehat{A}_{11}&\widehat{D}_{50}\\ &\widehat{D}_{61}&\widehat{A}_{20}&\widehat{A}_{21}&\widehat{B}_{30}&\widehat{B}_{31}&\widehat{D}_{60}\end{array}$
Here each column corresponds to a circuit time step, and each row to a block row of the matrices (67), (71), with secondary superscripts labeling the location corresponding to a free term (“0”) and the other monomial (“1”). Circuits corresponding to the original matrices with 2nd, 3rd, or 4th row removed were constructed identically, with the help of column permutations rendering the matrices in the form (67), (71), up to block labels.
Based on this addressing scheme for a single measurement round, we designed three types of four-round syndrome measurement circuits, $1$, 1212, and 1234555To guarantee rank preservation after removal of the 1st block row (which is important for the $1$ cycle), guided by Lemma 6, the polynomials were ordered so that $a_{4}=1+x$ [which gives block $D$ in Eqs. (45) and (52)].. Here a digit stands for the row omitted in each particular round of measurement, and we constructed detector events matching pairs of closest available measurement outcomes, either from neighboring or next-nearest-neighboring rounds. Overall, all our circuits have fixed duration of 9 measurement rounds, including two four-round cycles and a final measurement of all data qubits in $Z$ or $X$ basis, depending on the initial state of the data qubits, |$\ket{0}^{\otimes n}$ or |$\ket{+}^{\otimes n}$. The results of the final measurement are combined to construct the final set of detector events (thus computed syndrome bits are exact), and also the logical observables which allow to verify whether the decoding was successful.
Thus designed measurement protocols have been implemented as Stim circuits[53], using CX and XCX gates, with a standard circuit error model. Namely, we included probability-$p$ single-qubit depolarizing noise on the data qubits at the start of each round, two-qubit depolarizing noise with probability $p$ after every two-qubit gate, and also bit-flip (phase-flip) Pauli error channel with the same probability $p$ before and after each $Z$-basis ($X$-basis) measure-reset (MR) gate. There are no single-qubit gates in our circuits, and we chose to omit errors for idling intervals due to the fact that in most platforms single-qubit Clifford gates and idle intervals have error rates small compared to those of two-qubit entangling gates or single-qubit MR operations.
For accurate decoding, one needs to know probabilities of individual faults in the circuit and their effects on the detectors and the logical observables. This information for a given circuit is computed automatically by Stim, which also aggregates the faults producing the identical effect, and can be exported as a detector error model (DEM) file[53]. Information about aggregated faults in a DEM file can be regarded as columns of stabilizer-generator $H_{X}$ and logical $L_{X}$ matrices (with probabilities for each column also known) for an asymmetric quantum CSS code. In particular, the distance $d_{Z}$ of this code is identified as the circuit distance. We used vecdec program [54] to export the matrices from the DEM files associated with individual circuits, and dist_m4ri program to compute the corresponding distances (with the distances $d\leq 6$ computed exactly with a greedy enumeration algorithm, and upper bounds for larger distances using a random information set algorithm[55], exact with a high probability).
Circuit simulations have been done with Pauli-frame circuit simulator Stim [53] written by Craig Gidney, and the decoding using vecdec program [54]. Specifically, before attempting belief-propagation (BP) decoding, the program tries to match the syndrome data against a pre-computed list of small-weight error clusters; such a simple pre-decoder gives a substantial acceleration, especially at smaller error probabilities. If pre-decoding fails, the corresponding clusters are discarded, and a BP decoder with serial schedule, using both average and instantaneous logarithmic likelihood ratios (whichever converges first) is engaged, with up to 10 steps. Finally, if that fails, the logarithmic likelihood ratios (LLRs) are transmitted to the OSD-1 decoder. A sample of circuit simulation data for codes $[[42,6,4]]$, [66,6,6]][66,6,6]][ 66 , 6 , 6 ] ], and $[[96,6,8]]$ (see Tab. 1), with the standard error model described earlier and the “1212” syndrome-addressing cycle, is shown on a threshold-style plot in Fig. 1. To speed up the decoding, we simplified the circuit by dropping all “minority” detector events [^6] Unlike with the BP+OSD package[60, 61] whose performance may be degraded by additional detector events, see in Appendix A of Ref. 62, for the decoder we used, removing minority detector events increases logical error rates, e.g., from $0.20\%$ to $0.36\%$ at $p=0.1\%$ with the $[[42,6,4]]$ code and the “1212” circuits. Asymptotic slope at small $p$ consistent with $d/2$ (where $d$ is the code distance) is in agreement with the circuit distances computed independently. The (pseudo)threshold position (crossing point) is at $p_{c}\approx 1.1\%$, in approximate agreement with the circuit simulation results for similar codes in Ref. 34. We note that our value is (marginally) higher than the values for toric and surface codes under the same error models.
Figure 1: Logical error rate as a function of the circuit noise parameter $p$. Results for short even-distance codes $[[42,6,4]]$, [66,6,6]][66,6,6]][ 66 , 6 , 6 ] ], and $[[96,6,8]]$ (see Tab. 1) are shown. Inset shows the success probability $1-p_{L}$ as a function of $p$ in the vicinity of the crossing point(s) at $p\approx 1.1\%$
In addition, to verify the “single-shot” properties of the constructed codes, we have also tried sequential one-step sliding window decoding[18, 57, 23], with windows of different sizes. Details of our implementation are given in Ref. 25.
Some simulation results for the code $[[42,6,4]]$ are shown in Fig. 2, which shows the logical error rates $p_{L}$ as a function of the decoding window sizes. These results correspond to data points shown with open symbols, with three sets corresponding to different addressing cycles as indicated in the caption. While all cycles give nearly identical error rates with full-window ($T=9$) decoding, the “$1$” cycle (which discards most of the redundant stabilizer generators) shows the highest error rates with single-shot ($T=1$) and two-shot ($T=2$) decoding protocols. In comparison, the circuits designed with “1234” and “1212” cycles give nearly the same decoding accuracy with $T=2$ (two-shot) and $T=9$ (full-window) decoding. Note that only during the rounds labeled with “1” a full-rank set of stabilizer generators is addressed; this explains a slightly better accuracy of the “1212” measurement cycle compared to “1234”.
Figure 2: Logical error rate for the $[[42,6,4]]$ AMC code using full detector error model (both $X$ and $Z$ detector events) as a function of the decoding window size $T$; $T=9$ corresponds to full-block decoding, while the window sizes $T<9$ correspond to a sliding window decoding protocol, see text. The curves labeled in the caption as “$1$”, “1234”, and “1212” correspond to different measurement cycles as explained in the text. Single-shot decoding ($T=1$) is the least accurate as a single fault can sometimes lead to a decoding error. Two-shot decoding, $T=2$, with the “1234” and “1212” measurement cycles is nearly as accurate as full-block decoding, $T=9$. This is unlike the “$1$” measurement cycle where most of the redundant checks are discarded.
V Conclusion
In conclusion, we proposed an algebraic construction defining a family of CSS quantum error-correcting codes, whose main advantage is a highly redundant set of minimum-weight stabilizer generators in codes with relatively short block lengths. Just like the higher-dimensional QHP codes which our codes generalize, the construction defines a sequence of chain complexes of varying lengths. The constructed $D$-dimensional codes are highly-symmetric and they are characterized by confinement profiles and single-shot properties similar to those of the $D$-dimensional QHP codes.
It is instructive to compare the parameters of $D$-dimensional AMC codes over $F[C_{\ell}]$ with a characteristic polynomial $h(x)$, and $D$-dimensional QHP codes constructed from a linear cyclic code with parameters $[\ell,\kappa,\delta]$ and the same check polynomial $h(x)$. At level $j$, the QHP code has block length $N={D\choose j}\ell^{D}$, dimension $K={D\choose j}\kappa^{D}$, and minimal distance $D=\min(\delta^{j},\delta^{D-j})$, i.e., its rate scales as power of that of the original code, $R=(\kappa/\ell)^{D}$. In comparison, the block lengths and the dimensions of AMC codes scale linearly with $\ell$, $n={D\choose j}\ell$ and $k={D\choose j}\kappa$, so that the rate does not change, which is also a feature of GB codes.
While there is no explicit expression for the distance, we can use the codes in Tab. 1 to compare the actual distances with the corresponding upper bounds from Statement 5, which for these codes are given by the block size, $\ell$. For $d\leq 9$, we have a pretty good linear scaling, $d\geq\ell/2$, while for the remaining three codes, $d>\ell/3$. (Unfortunately, we do not expect the distances to remain as high for larger codes.)
Another nice property of the proposed construction is the high redundancy of minimum-weight stabilizer generators. While for a two-block code of dimension $k=2\kappa$ there are exactly $k$ minimum-weight redundant generators of each kind, stabilizer generator matrices of a balanced AMC code with $D=2j$ have ranks (${2j-1\choose j-1}(\ell-\kappa)$ and (${2j\choose j-1}\ell$ rows, which gives over a quarter redundant rows for $D=4$ and over a third for $D=6$. In particular, this redundancy can be used to speed up the decoding and improve its accuracy in a realistic setting of circuit noise. This is illustrated by our circuit simulations with relatively short AMC codes, which show two-shot decoding almost as accurate as full-block decoding, and a (pseudo)threshold above $1\%$, higher than that for surface codes.
Acknowledgements.
This work was supported in part by the APS M. Hildred Blewett Fellowship (HKL), and the NSF awards 2112848 (LPP) and OIA-2044049 (AAK).
Appendix A Proof of Statement 2
The calculation is based on the fact that the group algebra $F[C_{\ell}]$, isomorphic to the ring $R\equiv F[x]/(x^{\ell}-1)$ of modular polynomials, is a gcd domain. That is, for any two polynomials $a(x),c(x)\in F[x]/(x^{\ell}-1)$, there exists a polynomial $h_{ac}(x)\equiv\mathop{\rm gcd}(a,c)\in R$, the greatest common divisor of the two original polynomials. Moreover, there also exist Bézout coefficients $u(x),v(x)\in F[x]/(x^{\ell}-1)$ such that
$u(x)a(x)+v(x)c(x)=h_{ac}(x)\equiv\mathop{\rm gcd}(a,c).$ (73)
Equivalently, if we introduce the rescaled polynomials $a^{\prime}(x)=a(x)/h_{ac}(x)$, $c^{\prime}(x)=c(x)/h_{ac}(x)$, this gives $u(x)a^{\prime}(x)+v(x)c^{\prime}(x)=1$. Rescaled polynomials and the Bézout coefficients can be used to construct unimodular (and thus invertible) matrices,
$M_{\pm}(x)=\left(\begin{array}[c]{cc}a^{\prime}(x)&\pm c^{\prime}(x)\\ \mp v(x)&u(x)\end{array}\right),\quad\det M_{\pm}=1.$
The corresponding two-block matrices are also invertible,
$\hat{W}_{\pm}=\left(\begin{array}[c]{cc}A^{\prime}&\pm C^{\prime}\\ \mp V&U\end{array}\right),\quad\hat{W}_{\pm}^{-1}=\left(\begin{array}[c]{cc}U&\mp C^{\prime}\\ \pm V&A^{\prime}\end{array}\right),$
where $A^{\prime}=\mathop{\mathrm{M}}\nolimits(a^{\prime})$, $C^{\prime}=\mathop{\mathrm{M}}\nolimits(c^{\prime})$, etc. are $\ell\times\ell$ circulant matrices constructed from the corresponding polynomials; we have $AC^{\prime}-CA^{\prime}=0$ and $UA+VC=H_{ac}\equiv\mathop{\mathrm{M}}\nolimits(h_{ac})$.
Our approach is to construct a sequence of row and column transformations diagonalizing the boundary operator matrices in an AMC complex. First, demonstrate the steps on the first two matrices in the 3-complex ${\cal R}=\mathop{\text{\small\sc AMC}}\nolimits(a,b,c)$, see Eqs. (23) and (27). Here, we want to eliminate matrices $C=\mathop{\mathrm{M}}\nolimits(c)$, while replacing all matrices $A$ with $H_{ac}=\mathop{\mathrm{M}}\nolimits(h_{ac})$. To this end, denote $W^{\pm}_{i,j}$ the unimodular block-diagonal matrices constructed from an identity matrix by replacing blocks in positions $\{(i,i),(i,j),(j,i),(j,j)\}$ with the blocks of the corresponding matrices $\hat{W}_{\pm}$ (row-first order), $\overline{W}^{\pm}_{i,j}$ the corresponding inverse matrices, and $W^{\pm}_{i,j}$, $\overline{W}^{\pm}_{i,j}$ the same matrices with additional sign-flip. Explicitly,
$Q_{1}^{\prime}$ $=$ $Q_{1}\overline{W}^{+}_{31}=[C,B,A]\left[\begin{array}[c]{ccc}A^{\prime}&&V\\ &I\\ -C^{\prime}&&U\end{array}\right]=[0,B,H_{ac}],$
$Q_{2}^{\prime}$ $=$ ${W}^{+}_{3,1}Q_{2}\overline{W}^{-}_{3,1}$
$=$ $\left[\begin{array}[c]{ccc}+U&&-V\\ &I\\ +C^{\prime}&&+A^{\prime}\end{array}\right]\left[\begin{array}[c]{ccc}B&A&0\\ -C&0&A\\ 0&-C&-B\end{array}\right]\overline{W}^{-}_{3,1}$
$=$ $\left[\begin{array}[c]{cc|c}UB&H_{ac}&VB\\ \hline\cr-C&0&A\\ C^{\prime}B&0&-A^{\prime}B\end{array}\right]\left[\begin{array}[c]{ccc}A^{\prime}&&-V\\ &I\\ C^{\prime}&&U\end{array}\right]$
$=$ $\left[\begin{array}[c]{cc|c}B&H_{ac}&0\\ \hline\cr 0&0&H_{ac}\\ 0&0&-B\end{array}\right].$
The matrices are now reduced to a block-diagonal form, with the blocks being the matrices from the 2-dimensional complex $\mathop{\text{\small\sc AMC}}\nolimits(h_{ac},b)$, which, with the help of the identity
$\mathop{\rm gcd}(h_{ab},c,x^{\ell}-1)=\mathop{\rm gcd}(a,b,c,x^{\ell}-1)\equiv h(x),$
immediately gives
$\mathop{\rm rank}Q_{1}$ $=$ $\mathop{\rm rank}Q_{1}^{\prime}=\ell-\deg h=\ell-\kappa,$
$\mathop{\rm rank}Q_{2}$ $=$ $\mathop{\rm rank}Q_{2}^{\prime}=2(\ell-\deg h)=2(\ell-\kappa).$
The elimination process goes the same way for larger-$D$ complexes. Due to symmetry, whenever a pair of matrices, say, $A$ and $D$, are located in positions $(x_{1},y_{1})$, $(x_{1},y_{2})$ (same column) and in $(x_{2},y_{3})$, $(x_{3},y_{3})$ (same row), some other matrix, e.g., $B$, the $3\times 3$ block selected (rows $\{y_{1},y_{2},y_{3}\}$, columns $\{x_{1},x_{2},x_{3}\}$) will necessarily have the form of $Q_{2}$, up to some signs and the block labels.
We illustrate the elimination process on the example of the first two matrices from the 4-complex $\mathop{\text{\small\sc AMC}}\nolimits(a,b,c,d)$ in Eqs. (40), (45). Here $ua+vd=h_{ad}\equiv\mathop{\rm gcd}(a,d,x^{\ell}-1)$, $a^{\prime}=a/h_{ad}$, $d^{\prime}=d/h_{ad}$, and capital letters correspond to block matrices, $A^{\prime}=\mathop{\mathrm{M}}\nolimits(a^{\prime})$, $U=\mathop{\mathrm{M}}\nolimits(u)$, etc.
$R_{1}^{\prime}$ $=$ $R_{1}\overline{W}^{+}_{4,1}=[D,C,B,A]\left[\begin{array}[c]{cccc}A^{\prime}&&&V\\ &I\\ &&I\\ -D^{\prime}&&&U\end{array}\right]=[0,C,B,H_{1}],\text{ and}$
$R_{2}^{\prime}$ $=$ $W^{-}_{4,1}R_{2}\overline{W}^{-}_{5,1}\overline{W}^{-}_{6,2}$
$=$ $\left[\begin{array}[c]{cccc}U&&&-V\\ &I\\ &&I\\ D^{\prime}&&&A^{\prime}\end{array}\right]\left[\begin{array}[c]{ccc|ccc}C&B&A\\ \hline\cr-D&&&B&A&0\\ &-D&&-C&0&A\\ &&-D&0&-C&-B\\ \end{array}\right]\ldots$
$=$ $\left[\begin{array}[c]{ccc|ccc}UC&UB&H_{1}&0&VC&VB\\ \hline\cr-D&&&B&A&0\\ &-D&&-C&0&A\\ D^{\prime}C&D^{\prime}B&0&0&-A^{\prime}C&-A^{\prime}B\\ \end{array}\right]\left[\begin{array}[c]{cccccc}A^{\prime}&&&&-V\\ &I\\ &&I\\ &&&I\\ D^{\prime}&&&&U\\ &&&&&I\end{array}\right]\overline{W}_{6,2}^{-}$
$=$ $\left[\begin{array}[c]{ccc|ccc}C&UB&H_{1}&0&0&VB\\ \hline\cr 0&&&B&H_{1}&0\\ &-D&&-C&0&A\\ &D^{\prime}B&0&0&-C&-A^{\prime}B\\ \end{array}\right]\left[\begin{array}[c]{cccccc}I\\ &-V&&&&A^{\prime}\\ &&I\\ &&&I\\ &&&&I\\ &U&&&&D^{\prime}\end{array}\right]=\left[\begin{array}[c]{ccc|ccc}C&B&H_{1}&&&\\ \hline\cr 0&&&B&H_{1}&0\\ &0&&-C&0&H_{1}\\ &&0&0&-C&-B\\ \end{array}\right].$
The proof can be now completed recursively. Indeed, the rank expression in Statement 2 is valid for $D=2$ which corresponds to a GB code based on a cyclic group. Assuming it is valid for some $D\geq 2$, take a circulant matrix $N=\mathop{\mathrm{M}}\nolimits(n(x))$ and use Eqs. (33), (36), and ((39) to construct the boundary operators for the dimension $D+1$. Now, use the reduction protocol described to eliminate matrices $N$ and replace $A\to\mathop{\mathrm{M}}\nolimits(h_{a,n})$, where $h_{a,n}=\mathop{\rm gcd}(a,n,x^{\ell}-1)$. By assumption, the ranks at level $j$ are $\mathop{\rm rank}Q_{j}={D-1\choose j-1}(\ell-\kappa)$, where $\kappa=\deg h(x)$, and the expression for the characteristic polynomial now includes all $D+1$ polynomials including $n(x)$. This gives
$\mathop{\rm rank}R_{j}^{D+1}$ $=$ (${D-1\choose j-2}(\ell-\kappa)+{D-1\choose j-1}(\ell-\kappa)$
$=$ (${D\choose j-1}(\ell-\kappa),$
in agreement with the assumption. This completes the induction. ∎
Appendix B Proof of Statement 3
The proof is based on the following:
Lemma 6.
For a $D$-complex ${\cal Q}=\mathop{\text{\small\sc MBC}}\nolimits(A,B,\ldots)$, suppose one of the matrices, e.g., matrix $A$, which does not limit generality, can be factored out from the other matrices, that is, $B=B^{\prime}A=AB^{\prime}$, $C=C^{\prime}A=AC^{\prime}$, etc., then $\mathop{\rm rank}Q_{j}^{(D)}={D-1\choose j-1}\mathop{\rm rank}A$.
Proof.
Permute the terms in the product so that $A$ enters last, and use the equations (33), (36), and (39). The block $N=A$ enters $Q_{j}$ only along the diagonal of the square block of block-dimension (${D-1\choose j-1}$; the remaining matrices are above and to the right. Using the assumed factorization, $B=B^{\prime}A$, etc., row and column transformation can be used to diagonalize the matrix $Q_{j}$, leaving only the square block with $A$ along the diagonal. This gives $\mathop{\rm rank}Q_{j}={D-1\choose j-1}\mathop{\rm rank}A$. ∎
Proof of Statement 3. By assumption, the group algebra $F[G]$ is semisimple; every ideal in such a group algebra is generated by an idempotent element. That is, for any $a\in F[G]$, there is an idempotent element $e_{a}\in F[G]$ (necessarily $e_{a}=e_{a}^{2}$) such that $e_{a}a=a$ and, necessarily, $(1-e_{a})a=0$. Then, given $\{a,b,c,\ldots\}\subset F[G]$, construct the corresponding matrices $A=\mathop{\mathrm{M}}\nolimits(a)$, $B=\mathop{\mathrm{M}}\nolimits(b)$, etc, as well as the idempotent matrices $E_{a}=\mathop{\mathrm{M}}\nolimits(e_{a})$, $E_{b}=\mathop{\mathrm{M}}\nolimits(e_{b})$, etc; all of these matrices commute and satisfy, e.g., $E_{A}A=AE_{A}=A$, $(I-E_{A})A=0$. For a $D$-complex $\mathop{\text{\small\sc AMC}}\nolimits(a,b,c,\ldots)$, there are $D$ group algebra elements and $D$ corresponding idempotent matrices.
Further, by the same assumption, for any $a\in F[G]$, there exists a pseudoinverse $x\equiv\underline{a}\in F[G]$ such that $ax=e_{a}ax=e_{a}$. Denote the corresponding matrix $\underline{A}=\mathop{\mathrm{M}}\nolimits(\underline{a})$, with the property $A\underline{A}=E_{A}$. Together with $A=E_{A}A$, this implies
$\mathop{\rm rank}(E_{A})=\mathop{\rm rank}(A).$
Given a binary string $\boldsymbol{\alpha}=(\alpha_{1},\alpha_{2},\ldots,\alpha_{D})\in\mathbb{F}_{2}^{D}$, consider the idempotent matrices
$E_{\boldsymbol{\alpha}}\equiv\prod_{i=1}^{D}E_{i}^{\alpha_{i}}(I-E_{i})^{1-\alpha_{i}}.$
By construction, these commute with all block matrices $A$, $B$, …, and satisfy the following orthogonality and completeness properties,
$E_{\boldsymbol{\alpha}}E_{\boldsymbol{\alpha}^{\prime}}=\delta_{\boldsymbol{\alpha},\boldsymbol{\alpha}^{\prime}},\quad\sum_{\boldsymbol{\alpha}\in\mathbb{F}_{2}^{D}}E_{\boldsymbol{\alpha}}=I_{\ell}.$ (82)
Consider a boundary operator (matrix) $Q_{j}$, $1\leq j\leq D$ of a known block-size $r\times c$, and the corresponding matrix $Q_{j}(\boldsymbol{\alpha})$ with each block row multiplied by $E_{\boldsymbol{\alpha}}$. By properties of idempotent matrices $E_{\boldsymbol{\alpha}}$, non-zero rows in $Q_{j}(\boldsymbol{\alpha})$ with different $\boldsymbol{\alpha}$ are linearly independent. Also, this is an invertible transformation, thus
$\mathop{\rm rank}Q_{j}=\sum_{\boldsymbol{\alpha}\in\mathbb{F}_{2}^{D}}\mathop{\rm rank}Q_{j}(\boldsymbol{\alpha}).$
Now, we know that $Q_{j}(\boldsymbol{0})$ is identically zero. Consider $Q_{j}(\boldsymbol{\alpha})$ for $\boldsymbol{\alpha}\neq 0$. Without limiting generality, $\alpha_{1}\neq 0$. Due to the existence of the pseudoinverse $\underline{A}$, for every other (e.g., originating from the matrix $B$) we have
$E_{\boldsymbol{\alpha}}B=E_{\boldsymbol{\alpha}}E_{A}A\underline{A}B=E_{\boldsymbol{\alpha}}A\,\underline{A}B.$
This expression is valid even if $E_{\boldsymbol{\alpha}}A=0$, in which case $E_{\boldsymbol{\alpha}}=0$. Evidently, the projected blocks satisfy the conditions of the Lemma 6, which gives
$\mathop{\rm rank}Q_{j}(\boldsymbol{\alpha})={D-1\choose j-1}\mathop{\rm rank}E_{\boldsymbol{\alpha}}A={D-1\choose j-1}\mathop{\rm rank}E_{\boldsymbol{\alpha}},$
where $\alpha_{1}\neq 0$. The same construction works with other non-zero $\boldsymbol{\alpha}$, while for $\boldsymbol{\alpha}=\boldsymbol{0}$ the projected matrix is identically zero, and thus $\mathop{\rm rank}Q_{j}(\boldsymbol{0})=0$. This gives, finally, with the help of the completeness relation in Eq. (82):
$\mathop{\rm rank}Q_{j}$ $=$ $\sum_{\boldsymbol{\alpha}\in\mathbb{F}_{2}^{D}}\mathop{\rm rank}Q_{j}(\boldsymbol{\alpha})$
$=$ $\sum_{\boldsymbol{\alpha}}(1-\delta_{\boldsymbol{0},\boldsymbol{\alpha}}){D-1\choose j-1}\mathop{\rm rank}E_{\boldsymbol{\alpha}}$
$=$ (${D-1\choose j-1}[\mathop{\rm rank}I_{\ell}-\mathop{\rm rank}E_{\boldsymbol{0}}]$
$=$ (${D-1\choose j-1}(\ell-\kappa),\quad\kappa\equiv\mathop{\rm rank}E_{\boldsymbol{0}},$
with $\kappa$ being the rank of the products of idempotent matrices $(I-E_{A})$, $(I-E_{B})$, …, complementary to the commuting blocks $A$, $B$, …used to construct the $D$-complex.
The proof is completed by noticing that for $j=D$, the binomial coefficient equals to $1$, thus $\mathop{\rm rank}Q_{D}=\ell-\kappa$.∎
Appendix C Proof of Statement 4
The proof amounts to an observation that the construction gives a $D$-dimensional QHP code constructed from the $F$-linear abelian group code with the check matrix $A=\mathop{\mathrm{M}}\nolimits(a)$. With $k=0$, all homology groups are trivial, and all distances are infinite. Otherwise, the $1$-complex with the boundary operator $A$ has both subspaces of dimensions $\ell$, both homology groups of dimensions $k$, and the homological distances $d_{0}^{(1)}=1$, $d_{1}^{(1)}=d$. This gives the parameters of the product complex: homology group dimensions $k_{j}$ follow immediately the Künneth formula (13) as the coefficient of $x^{j}$ in the Poincaré polynomial $(k+xk)^{D}$, subspace dimensions similarly from Eq. (12) written as a Poincaré polynomial $(\ell+x\ell)^{D}$, and distances from Theorem 17 in Ref. 21.
Appendix D Proof of Statement 5
It is trivial to verify that the Statement is true for $j=0$ and for $j=D$. Denote $r\equiv m_{j}^{D}={D\choose j}$ and $c\equiv m_{j+1}^{D}={D\choose j+1}$ be the block dimensions of subspaces ${\cal Q}_{j}$ and ${\cal Q}_{j+1}$, respectively. For any non-zero vector $\boldsymbol{u}$ in $C_{h}^{\perp}$, we can construct $r$ same-weight vectors $\boldsymbol{u}_{i}$ in the space ${\cal Q}_{j}$ (with non-zero elements in only one block, and that equal to $\boldsymbol{x}$) which satisfy the equation $Q_{j}\boldsymbol{u}_{i}=\boldsymbol{0}$. We would like to show that all of them cannot be homologically trivial at the same time.
To this end, define $g(x)=(x^{\ell}-1)/h(x)$ and write the non-zero codeword as $u(x)=i(x)g(x)$, with $\deg i(x)<k$, and assume all $r$ vectors $\boldsymbol{u}_{i}$ to be trivial. This can be written as the matrix-polynomial equation
$u(x)I_{r\times r}=Q_{j}(x)\Xi_{r\times c}(x)\bmod x^{\ell}-1,$
where $\Xi_{r\times c}(x)$ is a matrix of specified dimensions over $F[x]/(x^{\ell}-1)$ which solves the equation. With the help of the invertible transformations similar to those constructed in Sec. A, matrix $Q_{j}(x)$ can be transformed to a diagonal form,
$Q_{j}(x)=W_{L}(x)h(x)\Lambda_{r\times c}W_{R}(x),$
with square matrices $W_{L}(x)$ and $W_{R}(x)$ invertible in the ring $F[x]/(x^{\ell}-1)$, and $\Lambda_{r\times c}$ with only
$\mathop{\rm rank}\Lambda_{r\times c}=m_{j}^{D-1}<\min(r,c)$
non-zero elements (equal to 1), all in different rows and columns. Using the fact that the object on the left-hand-side (l.h.s.) is proportional to the identity matrix, we have
$u(x)I_{r\times r}$ $=$ $W_{L}(x)h(x)\Lambda_{r\times c}W_{R}(x)\,\Xi_{r\times c}(x)\bmod x^{\ell}-1,$
$u(x)I_{r\times r}$ $=$ $h(x)\Lambda_{r\times c}{\Gamma_{r\times c}(x)}\bmod x^{\ell}-1,$
where the coefficients in $\Xi(x)$ have been replaced with the coefficients in $\Gamma(x)\equiv W_{R}(x)\,\Xi_{r\times c}(x)\,W_{L}(x)$; this transformation is one-to-one since the transformation matrices are invertible by construction.
If we pick a zero column in $\Lambda_{r\times c}$, the equation reads $u(x)=0\bmod x^{\ell}-1$, contradicting the initial assumption that $u(x)\neq 0$. Thus, for any non-zero codeword in $C_{h}^{\perp}$, there is a non-trivial vector in $H_{j}(\mathcal{Q})$ of the same weight, which proves the stated upper bound on the distance.∎